# DND: Boosting Large Language Models with Dynamic Nested Depth

### v1

@@ 5 Conclusion
In this work, we introduce Dynamic Nested Depth (DND), a novel and efficient method for enhancing Large Language Model performance. DND adaptively identifies critical tokens and selectively deepens their computation via nested re-processing. This is achieved through a token-choice routing design with a normalized output fusion strategy. The precision and stability of this selection process are guaranteed by our router controlling loss and threshold control scheme. We validated DND on both dense (Qwen3-1.7B) and larger-scale sparse MoE (Qwen3-30B-A3B) models, demonstrating substantial accuracy improvements with a negligible parameter increase (<< 0.1M) and minimal computing increase. These results affirm that targeted, dynamic nested depth computation is a powerful method for boosting LLMs’ performance.
@@ Appendix B More Details
The post-training stage uses the AdamW optimizer (Loshchilov & Hutter, 2017), with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.95, weight decay of 0.1, and gradient clipping at 1.0. During the post-training stage, we employ a cosine learning rate scheduler with a learning rate of $5\times 10^{-6}$ that gradually decays to a minimum of $1\times 10^{-6}$. All experiments are run on H100 GPUs. For Qwen3-1.7B, training is conducted for two epochs across 128 GPUs, taking one day. For Qwen3-30B-A3B, training is conducted for four epochs across 256 GPUs, taking approximately three days.
@@ Appendix C Detailed Comparisons with MOR
Model Scale Limited to models with around 1B parameters. Applied from 1.7B dense to 30B MoE models.
@@
Appendix D More Visualizations
@@
Appendix E Limitations and Future Works

### v2

@@ 4.5 Token Selection Analysis
Why critical tokens are selected? As shown in Fig. 4(a), to examine whether the tokens selected by our DND model are indeed critical, we analyze the relation between the selection frequency of tokens routed by Qwen3-30B-A3B+DND and the vanilla model’s logit entropy without any reprocessing. According to the findings in (entropy), high entropy in the logit indicates that the model is uncertain about which vocabulary to select or is hesitating between multiple possible answers. We find that tokens with higher logit entropy are frequently selected by the router across multiple layers. This shows that DND preferentially selects tokens with greater uncertainty, validating the motivation behind critical token selection and confirming the effectiveness of the router.
Why better representations are learned? Furthermore, as shown in Fig. 4(b), to validate that our DND strategy reduces the model’s hesitation or uncertainty about critical tokens, we evaluated the variation in logit entropy for the same token after applying the DND strategy, compared to the entropy in the vanilla model. We found that after nested reviews, the logit entropy of the selected tokens significantly decreased, proving the effectiveness of our method.
@@ 5 Conclusion
In this work, we introduce Dynamic Nested Depth (DND), a novel and efficient method for enhancing Large Language Model performance. DND adaptively identifies critical tokens and selectively deepens their computation via nested re-processing. This is achieved through a token-choice routing design with a normalized output fusion strategy. The precision and stability of this selection process are guaranteed by our router controlling loss and threshold control scheme. We validated DND on both dense (Qwen3-1.7B, Llama3.2-1B, Gemma3-1B) and larger-scale sparse MoE (Qwen3-30B-A3B) models, demonstrating substantial accuracy improvements with a negligible parameter increase (<< 0.1M) and minimal computing increase. These results affirm that targeted, dynamic nested depth computation is a powerful method for boosting LLMs’ performance.
@@ Appendix B More Details
The post-training stage uses the AdamW optimizer (Loshchilov2017DecoupledWD), with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.95, weight decay of 0.1, and gradient clipping at 1.0. During the post-training stage, we employ a cosine learning rate scheduler with a learning rate of $5\times 10^{-6}$ that gradually decays to a minimum of $1\times 10^{-6}$. All experiments are run on H100 GPUs. For Qwen3-1.7B, training is conducted for two epochs across 128 GPUs, taking one day. For Qwen3-30B-A3B, training is conducted for four epochs across 256 GPUs, taking approximately three days. For Llama-3.2-1B and Gemma3-1B, training is conducted for two epochs across 64 GPUs, taking approximately 1-2 days.
@@ Appendix C Detailed Comparisons with MOR
Model Scale Limited to models with around 1B parameters. Applied from 1B dense to 30B MoE models.
Purpose
Achieve comparable performance with fewer
parameters by using dynamically cycled depth
under the same FLOPs budget.
Achieve better performance with almost unchanged
parameter count by adding only a small amount of
extra depth computation.
@@
Appendix D More Token Selection Analysis
Figure 8: Analysis of the Relationship between DND’s Selection Preference and the Magnitude of Hidden State Changes across Transformer Layers. DND exhibits a stronger preference for selecting tokens whose representations undergo larger changes after passing through a given layer.
Beyond the entropy analysis of token logits in Sec. 4.5, we further examined another selection pattern of DND across layers. Specifically, we measured the variation in each token’s hidden state across transformer layers. As shown in Fig. 8, we found that DND tends to select tokens whose representations change more strongly after passing through a certain transformer layer. This suggests that the model may be less confident about such tokens and therefore chooses to revisit them to obtain a more reliable representation.
Appendix E More Visualizations
@@
Appendix F Limitations and Future Works